OpenStudy (anonymous):

find the center and radius of the circle given the circle x^2+y^2-18x+45=0 is tangent to the line y=(4/3)x-2

OpenStudy (anonymous):

complete the square. \[x^2-18x +y^2+45 =0\] \[x^2-18x +81 +y^2 =-45+81\] \[(x-9)^2 +y^2 =-45+9\] \[(x-9)^2 +(y+0)^2=-36\] Center: (9,0) Radius: \[sqrt{36}\] =6

OpenStudy (anonymous):

o I did that and got the same thing but was wondering y they give the tangent line y=4/3x-2 ?? we don't have anything t do with that line??

OpenStudy (campbell_st):

seems to be a calculation error is you add 81 to complete the square in x, then you need to do it on both sides of the equation.... that's causing the problem with finding the radius \[(x -9)^2 + y^2 = -45 + 81\]

OpenStudy (campbell_st):

so when you have the centre and radius... so there are several steps to the next part... 1. substitute the equation of the line into the circle equation and find how many points of intersection. 2. find the distance from the centre to the line to see if it is equal to the radius... given the way the question is written I'd say option 2. so if you have a line in standard form \[Ax + By + C = 0\] the distance from a point \[(x_{0}, y_{0})\] can be found using \[d = \frac{\left| Ax_{0} + By_{0} + C \right|}{\sqrt{A^2 + B^2}}\] hope it helps...

OpenStudy (anonymous):

so\[\frac{ 4 }{ 3 }(9)+1(0)-2/\sqrt{(16/9+1)}=6\]

OpenStudy (anonymous):

that equals to 6... which is the radius ...so what does this show ??

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I just want to be clear on this @campbell_st

OpenStudy (anonymous):

@amistre64 what does it show... can u say??

OpenStudy (amistre64):

i wouldnt know why the tangent is given ... the equation of a circle is defined already by its center and radius of you complete the squares.

OpenStudy (campbell_st):

well if the tangent is y = 4/3x - 2 then 3y = 4x - 6 or 4x - 3y - 6 = 0 so you need the centre of the circle (9, 0) and the radius 6 so all you are going to do is show that the distance from the tangent to the radius is 6 units \[d = \frac{\left| 4 \times 9 - 3 \times 0 - 6 \right|}{\sqrt{4^2 + (-3)^2}}\] so this results in \[d = \frac{30}{5} \] so the distance from the centre to the tangent is 6... so you have the correct equation by completing the square...

OpenStudy (campbell_st):

oops should read ... show the distance from the centre to the tangent is 6 units

OpenStudy (anonymous):

ok so the distance and the radius is the same... does this always happen..just curious ..i have a test 2morrow

OpenStudy (amistre64):

i do not read in the post that we are to show the distance from the center to the tangent line is required ...

OpenStudy (amistre64):

find radius and center, given _____ otherwise it would read, find and show that tangent line is yada yada .... at least to me that is

OpenStudy (campbell_st):

well its a silly question... but if you thing about it, the perpendicular distance from the centre to a tangent will always be a radius

OpenStudy (campbell_st):

I meant the initial question is silly

OpenStudy (campbell_st):

a tangent has 1 point of contact on the circumference... so the distance and radius will be the same.

OpenStudy (anonymous):

yea i know i hope no question like this come for this test that i have tomorrow ...

OpenStudy (amistre64):

just make sure you read the test questions, and determine what it is that they are asking for. once you have answered them, all else is extraneous :)

OpenStudy (anonymous):

ok will do that thanks